# Filling the gap in the IERS C01 polar motion series in 1858.9–1860.9

### v1

@@ 4 SSA analysis
For the C01 time series, we will consider a model in the form of signal plus noise, where the signal consists of three main components: trend, AW, and CW, with possible addition of other smaller components. Unlike the deterministic BAC model described above, the SSA analysis does not imply a priori restrictions on the trend, AW and CW variations.
We will process time series with 10 measurements per year. Thus, the AW component corresponds to a period of 10 points, and the CW component corresponds to a period of about 11.9 points. For further presentation of our method, we denote IERS C01 PM coordinates $X_{p}$ and $Y_{p}$ as $\mathsf{X}$ and $\mathsf{Y}$, respectively.
@@
Let us first describe the general approach to gap filling. Suppose there is a method that knows how to remove noise from a noisy signal. Then we can propose a simple iterative gap-filling method. First, the places with missing values are filled with some initial values, e.g., the mean of the series, then the signal extraction is performed, the obtained signal estimates are put at the gaps, the signal extraction method is applied again, and so the procedure is repeated until convergence.
@@
The SSA method in its version for signal extraction has only two parameters, the window length $L$ and the number of components $r$. At the first stage, the time series $\mathsf{X}=(x_{1},\ldots,x_{N})$ is transformed into the so-called trajectory matrix
@@
using the parameter $L$. Next, the singular value decomposition (SVD) of the trajectory matrix is constructed. Let $\mathbf{S}=\mathbf{X}\mathbf{X}^{\mathrm{T}}$, $\lambda_{1},\ldots,\lambda_{L}$ be the eigenvalues of the matrix $\mathbf{S}$ taken in non-decreasing order, $U_{1},\ldots,U_{L}$ be the orthonormalized system of eigenvectors corresponding to these eigenvalues. Let $d=\max\{k:\lambda_{k}>0\}$ and $V_{k}=\mathbf{X}^{\mathrm{T}}U_{k}/\sqrt{\lambda_{k}}$, $k=1,\ldots,d$. Then the SVD of the matrix $\mathbf{X}$ can be represented as a sum of elementary matrices
@@
In the second step, the first $r$ of the elementary matrices of the SVD are summed:
@@
and their sum is converted into a time series.
@@
It was shown in Golyandina et al. (2015) that if several signals have the same periods (with the same modulation), the Multivariate SSA (MSSA), a generalization of SSA for time series system analysis, extracts the signal more accurately. Such a method allows us to process $\mathsf{X}$ and $\mathsf{Y}$ series together.
@@
Taking into account the physical nature of the pole coordinates, this looks reasonable to consider a representation of the PM series as a complex time series. If the input signal consists of two harmonics with a phase shift of $\pi/2$, then the generalization of SSA for the analysis of complex time series extracts the signal even more accurately (Golyandina et al., 2015). In the case of Complex SSA (CSSA) for a time series of the form $x_{n}+{\mathrm{i}\mkern 1.0mu}y_{n},\ n=1,\ldots,N$, the transpose operator $\mathrm{T}$ is replaced by the Hermite conjugate operator $\mathrm{H}$. The complex form of SSA has been considered since its origin. In the complex form, the first descriptions of the method can be found, for example, in Kumaresan and Tufts (1982); Keppenne and Lall (1996).
To ensure more reliable results, we considered the three methods described above for reconstruction of PM signal: basic 1D-SSA for the series $\mathsf{Y}$, MSSA for the series ($\mathsf{X},\mathsf{Y}$), and CSSA for the series $\mathsf{X}+{\mathrm{i}\mkern 1.0mu}\mathsf{Y}$.
An important option for a practical SSA realization is a choice of the window length. Recommendations for optimal window selection include proportionality to the periods of the harmonics we want to extract (Golyandina et al., 2001). Therefore, for the window length $L$ we will consider two values, 119 and 238 points, since the AW and CW periods (10 and about 11.9, respectively) are included near integer number of times in windows of this length.
The parameter that we want to optimize using artificial gaps is $r$, the number of the SVD components assigned to the signal. We will call $r$ the model order. To find the optimal value of the model order, we vary it from 1 to 17 and calculate the error of gap filling according to Algorithm 1. We used the root-mean-square error (RMSE) and the mean absolute error (MAE) as a measure of the difference between the imputed values and the input time series values.
The mean MAE and RMSE errors for these models are presented in Table 2. These test results have shown that the most precise filling of artificial gaps were achieved using CSSA with $L$=119 and $r$ from 6 to 9. Similar results were obtained for RMSE, for which models of the same orders showed somewhat larger errors overall, but the optimal parameters $r$ and $L$ turned out to be the same as for the MAE error estimate. Thus, it has been found that the obtained results do not depend on the choice of the method for estimating the modeling error.
Table 2: MAE/RMSE filling errors (arcseconds) on artificial gaps for SSA model order $r=1\ldots 17$.
$r$ $L$=119 $L$=238
1D-SSA MSSA CSSA 1D-SSA MSSA CSSA
1 0.121/0.143 0.114/0.142 0.104/0.130 0.121/0.143 0.114/0.142 0.106/0.135
2 0.104/0.129 0.109/0.137 0.084/0.110 0.107/0.134 0.108/0.137 0.087/0.115
3 0.089/0.116 0.094/0.118 0.076/0.099 0.091/0.120 0.091/0.118 0.080/0.105
4 0.082/0.109 0.090/0.113 0.072/0.094 0.085/0.113 0.086/0.112 0.076/0.099
5 0.085/0.110 0.092/0.114 0.070/0.090 0.086/0.112 0.086/0.111 0.072/0.094
6 0.083/0.108 0.095/0.117 0.065/0.084 0.083/0.106 0.086/0.109 0.071/0.091
7 0.085/0.111 0.101/0.123 0.064/0.084 0.084/0.106 0.087/0.109 0.066/0.085
8 0.086/0.112 0.106/0.129 0.064/0.084 0.083/0.106 0.090/0.112 0.073/0.093
9 0.090/0.117 0.113/0.137 0.065/0.085 0.084/0.109 0.096/0.118 0.075/0.096
10 0.088/0.113 0.116/0.139 0.068/0.090 0.085/0.109 0.100/0.122 0.076/0.097
11 0.088/0.111 0.118/0.143 0.071/0.092 0.085/0.110 0.104/0.126 0.074/0.094
12 0.080/0.102 0.123/0.149 0.072/0.095 0.083/0.107 0.105/0.127 0.074/0.095
13 0.081/0.106 0.129/0.156 0.073/0.097 0.084/0.108 0.108/0.130 0.076/0.098
14 0.085/0.111 0.133/0.160 0.073/0.099 0.085/0.108 0.110/0.133 0.076/0.099
15 0.090/0.118 0.137/0.166 0.071/0.097 0.090/0.114 0.114/0.137 0.070/0.091
16 0.093/0.121 0.140/0.169 0.075/0.102 0.086/0.110 0.116/0.139 0.070/0.091
17 0.092/0.122 0.143/0.173 0.081/0.108 0.088/0.115 0.119/0.142 0.070/0.091
The gap-filling plots with the best parameters (CSSA, $L$=119, $r=6\ldots 9$) are shown in Fig. 4.
Figure 4: IERS C01 $Y_{p}$ series with filled missed data between 1858.9 and 1860.9 computed using the CSSA model with $L$=119 and $r$ from 6 to 9.
Consider the CSSA decomposition components of the time series $\mathsf{Y}$ after filling the actual gap. Recall that the imaginary parts of the complex components are related to the series $\mathsf{Y}$. The latter for the CSSA 6-component solution with $L=119$ are shown in Fig. 5, and Fig. 6 shows the $\mathsf{Y}$ components for the 9-component CSSA solution.
Figure 5: The $\mathsf{Y}$ components of CSSA solution with $r$=6 (in arcseconds) and their normalized spectra. The period of the most powerful oscillation and the contribution of the component to the total signal are shown on the right.
Figure 6: The $\mathsf{Y}$ components of CSSA solution with $r$=9 (in arcseconds) and their normalized spectra. The period of the most powerful oscillation and the contribution of the component to the total signal are shown on the right.
It can be seen the leading six components of the 9-component solution have practically the same periods as those for the 6-component solution. Some components show two spectral peaks, which can be caused by mixing of two periodicities or by phase instability of the component. In all solutions, the first component corresponds to the main CW oscillation, and the third component corresponds to the main AW oscillation.
Since the imputed values for the model orders $r$ from 6 to 9 (CSSA with $L=119$) provide approximately equal errors on the artificial gaps, the average of these values can be recommended as the final method of filling in the actual gap in the IERS C01 series. This filling method has a slightly smaller error (0.063 vs. 0.064–0.065) and a bit smaller error variation (standard deviation) (0.0251 vs 0.0252–0.0265) on the considered set of artificial gaps, in comparison with gap filling by $r$ components separately.
5 Final results
Average results (filling $Y_{p}$ values) computed by the BCA and SSA methods are presented in Table 3 with the precision corresponding to the IERS C01 series. The fourth column of Table 3 contains the differences between the results obtained by two methods, and the last column contains the average of the results obtained by two methods. The IERS C01 series with added filling values is shown in Fig. 7.
Table 3: Filled $Y_{p}$ values, arcseconds.
Epoch Model Difference Average
BCA SSA BCA–SSA
1858.9 $-0.24731$ $-0.24542$ $-0.00189$ $-0.24636$
1859.0 $-0.19886$ $-0.15757$ $-0.04129$ $-0.17822$
1859.1 $-0.11187$ $-0.04478$ $-0.06708$ $-0.07833$
1859.2 $-0.01403$ $0.05086$ $-0.06489$ $0.01841$
1859.3 $0.06414$ $0.09691$ $-0.03278$ $0.08053$
1859.4 $0.09921$ $0.08292$ $0.01628$ $0.09106$
1859.5 $0.08210$ $0.02188$ $0.06022$ $0.05199$
1859.6 $0.02039$ $-0.05820$ $0.07859$ $-0.01890$
1859.7 $-0.06490$ $-0.12716$ $0.06226$ $-0.09603$
1859.8 $-0.14678$ $-0.16397$ $0.01719$ $-0.15537$
1859.9 $-0.20112$ $-0.16264$ $-0.03847$ $-0.18188$
1860.0 $-0.21406$ $-0.13195$ $-0.08211$ $-0.17300$
1860.1 $-0.18546$ $-0.09002$ $-0.09543$ $-0.13774$
1860.2 $-0.12761$ $-0.05632$ $-0.07129$ $-0.09197$
1860.3 $-0.06015$ $-0.04340$ $-0.01675$ $-0.05178$
1860.4 $-0.00312$ $-0.05167$ $0.04855$ $-0.02740$
1860.5 $0.02922$ $-0.06956$ $0.09878$ $-0.02017$
1860.6 $0.03183$ $-0.07982$ $0.11165$ $-0.02399$
1860.7 $0.00889$ $-0.06961$ $0.07851$ $-0.03036$
1860.8 $-0.02926$ $-0.03929$ $0.01003$ $-0.03427$
1860.9 $-0.07053$ $-0.00424$ $-0.06628$ $-0.03739$
A better agreement between the two methods can be seen in the first year of the gap interval, 1859–1860. The agreement in the second year of the interval, 1860–1861, is somewhat worse, but can be considered satisfactory, considering that the $Y_{p}$ error in the IERS C01 series is 0.09′′. Generally speaking, it should be kept in mind that the accuracy of the reconstruction is limited by the accuracy of the IERS C01 data, especially before the gap in the 1840s and 1850s.
6 Conclusions
In this paper, the first attempt was made to address the problem of filling the 2-year gap in the IERS C01 $Y_{p}$ series in 1858.9–1860.9. Two methods were proposed and investigated for this purpose. The first is LS adjustment of the BCA model consisting of a bias and CW and AW harmonics with linearly changing amplitude. In a general case, such a model can be extended to more complicated trend and amplitude variations. The second is SSA for complex PM data, or CSSA, which allows us to get a fully data-driven solution not limited to the parameters of the deterministic BCA model.
We consider the result obtained by SSA as more reliable because it takes into account all the features of the observed PM data and all the main components of its structure, while the deterministic model uses only predefined components, which may not adequately describe the results of the observations. However, we present both the results obtained by the BCA and SSA methods in Table 3 so that the user can choose the variant that looks most appropriate for the problem under investigation, or simply use the average of the two.
Incorporating the proposed fill values into the IERS C01 series would provide continuous, evenly spaced series over the entire date range. This will allow for faster and more accurate analysis methods for PM analysis and mitigate a bias in PM modeling caused by the uneven spacing of the original IERS C01 series.

### v2

@@ 4 SSA analysis
For further presentation of our method, we denote IERS C01 PM coordinates $X_{p}$ and $Y_{p}$ as $\mathsf{X}$ and $\mathsf{Y}$, respectively. As in the case of the parametric model considered in Section 3, we will consider a model in the form of signal plus noise, assuming that the signal consists of three main components: trend, AW, and CW, with possible addition of other smaller components. Unlike parametric model specification, in the SSA method, the model does not need to be specified in advance, including setting the component periods. The signal components are defined adaptively by the SSA method and, as will be shown further on, are indeed composed of the listed components. We will process the PM time series with 10 measurements per year. Thus, the AW component corresponds to a period of 10 points, and the CW component corresponds to a period of about 11.9 points. Unlike the parametric BCA model described above, the SSA analysis does not imply a priori restrictions on the trend, AW and CW variations. The general approach to gap filling with SSA is described in Appendix A.
It was shown in Golyandina et al. (2015) that if several signals have the same periods (with the same modulation), Multivariate SSA (MSSA), a generalization of SSA for time series system analysis, extracts the signal more accurately. Such a method allows us to process $\mathsf{X}$ and $\mathsf{Y}$ series together.
Taking into account the physical nature of the pole coordinates, this looks reasonable to consider a representation of the PM series as a complex-valued time series and applied Complex SSA (CSSA). If the input signal consists of two harmonics with a phase shift of $\pi/2$, then CSSA extracts the signal even more accurately Golyandina et al. (2015).
To ensure more reliable results, we considered the three methods described above for reconstruction of PM signal: basic 1D-SSA for the series $\mathsf{Y}$, MSSA for the series ($\mathsf{X},\mathsf{Y}$), and CSSA for the series $\mathsf{X}+{\mathrm{i}\mkern 1.0mu}\mathsf{Y}$. Note that MSSA and CSSA are applied to both time series, $\mathsf{X}$ and $\mathsf{Y}$, together, so they take into account the relationships between the series. The methods MSSA and CSSA result in decompositions of both series, but we will use only the part of the decomposition of the series Y.
An important option for a practical SSA realization is a choice of the window length. Recommendations for optimal window selection include proportionality to the periods of the harmonics we want to extract (Golyandina et al., 2001). Therefore, for the window length $L$ we will consider three values, 60, 119 and 238 points, since the AW and CW periods, 1.0 and 1.19 yr, respectively, are included near integer number of times in windows of this length.
The parameter that we want to optimize using artificial gaps is $r$, the number of the SVD components assigned to the signal. We will call $r$ the model order. To find the optimal value of the model order, we vary it from 1 to 17 and calculate the error of gap filling according to Algorithm 1. We again used RMSE and MAE as a measure of the difference between the imputed values and the input time series values.
The mean MAE and RMSE errors for these models are presented in Table 3. These test results have shown that the most precise filling of artificial gaps were achieved using CSSA with ($L$=60, $r$=4, 5), ($L$=119, $r$=6–9), ($L$=238, $r$=7). These seven variants provide practically the same error of the approximation of the $\mathsf{Y}$ series. Similar results were obtained for RMSE, for which the optimal parameters $r$ and $L$ turned out to be the same as for the MAE estimate. Thus, it has been found that the obtained conclusions do not depend on the choice of the method for estimating the modeling error.
Table 3: MAE/RMSE filling errors on artificial gaps for SSA model order $r=1\ldots 17$, arcseconds.
$r$ $L$=60 $L$=119 $L$=238
1D-SSA MSSA CSSA 1D-SSA MSSA CSSA 1D-SSA MSSA CSSA
1 .125/.150 .117/.145 .104/.129 .121/.143 .114/.142 .104/.130 .121/.143 .114/.142 .106/.135
2 .112/.138 .114/.142 .084/.109 .104/.129 .109/.137 .084/.110 .107/.134 .108/.137 .087/.115
3 .093/.118 .101/.124 .070/.092 .089/.116 .094/.118 .076/.099 .091/.120 .091/.118 .080/.105
4 .081/.107 .102/.124 .066/.085 .082/.109 .090/.113 .072/.094 .085/.113 .086/.112 .076/.099
5 .080/.107 .110/.134 .064/.083 .085/.110 .092/.114 .070/.090 .086/.112 .086/.111 .072/.094
6 .078/.105 .116/.140 .067/.087 .083/.108 .095/.117 .065/.084 .083/.106 .086/.109 .071/.091
7 .081/.109 .124/.150 .067/.088 .085/.111 .101/.123 .064/.084 .084/.106 .087/.109 .066/.085
8 .078/.103 .131/.158 .071/.091 .086/.112 .106/.129 .064/.084 .083/.106 .090/.112 .073/.093
9 .083/.107 .139/.168 .073/.095 .090/.117 .113/.137 .065/.085 .084/.109 .096/.118 .075/.096
10 .086/.111 .144/.174 .077/.099 .088/.113 .116/.139 .068/.090 .085/.109 .100/.122 .076/.097
11 .094/.121 .148/.179 .079/.102 .088/.111 .118/.143 .071/.092 .085/.110 .104/.126 .074/.094
12 .092/.119 .150/.182 .083/.107 .080/.102 .123/.149 .072/.095 .083/.107 .105/.127 .074/.095
13 .095/.124 .152/.184 .085/.110 .081/.106 .129/.156 .073/.097 .084/.108 .108/.130 .076/.098
14 .097/.126 .153/.186 .087/.112 .085/.111 .133/.160 .073/.099 .085/.108 .110/.133 .076/.099
15 .103/.133 .155/.188 .087/.110 .090/.118 .137/.166 .071/.097 .090/.114 .114/.137 .070/.091
16 .105/.136 .156/.189 .085/.107 .093/.121 .140/.169 .075/.102 .086/.110 .116/.139 .070/.091
17 .110/.143 .157/.191 .094/.116 .092/.122 .143/.173 .081/.108 .088/.115 .119/.142 .070/.091
Comparison of the data presented in Table 2 and Table 3 shows that the accuracy of the $Y_{p}$ series approximation using CSSA method is a bit better than that for the BCA method. Remarkably, that the best RMSE in filling $\mathsf{Y}$ values achieved by both methods (0.08′′–0.09′′) are very close to the $Y_{p}$ uncertainty given in the IERS C01 series (0.090′′, which is assumed to be the standard deviation, equivalent to RMSE), confirming both the reality of the IERS C01 errors and the correctness of our error estimate. The gap-filling plots with the best parameters described above are shown in Fig. 4.
Figure 4: IERS C01 $Y_{p}$ series with filled missed data between 1858.9 and 1860.9 computed using the CSSA model with the seven models providing the minimal approximation error.
Consider the CSSA decomposition components of the time series $\mathsf{Y}$ after filling the actual gap. Recall that the imaginary parts of the complex components are related to the series $\mathsf{Y}$. The latter for the CSSA 6-component solution with $L=119$ are shown in Fig. 5, and Fig. 6 shows the $\mathsf{Y}$ components for the 9-component CSSA solution, as examples.
Figure 5: The $\mathsf{Y}$ components of CSSA solution with $r$=6 (in arcseconds) and their Lomb-Scargle periodograms (except the trend component, arbitrary units). The period of the most powerful oscillation and the contribution of the component to the total signal are shown on the right.
Figure 6: The $\mathsf{Y}$ components of CSSA solution with $r$=9 (in arcseconds) and their Lomb-Scargle periodograms (except the trend component, arbitrary units). The period of the most powerful oscillation and the contribution of the component to the total signal are shown on the right.
It can be seen that the leading six components of the 9-component solution have practically the same periods as those for the 6-component solution. Some components show two spectral peaks, which can be caused by mixing of two periodicities or by phase instability of the component. In all solutions, the first component corresponds to the main CW oscillation, and the third component corresponds to the main AW oscillation. In Figs. 5 and 6, two minor components with periods of 0.847 yr and 0.944 yr can also be seen. The corresponding peaks in Fig. 2 are small, at the noise level. We consider this fact as confirmation of the better sensitivity of the SSA method to small components in the analyzed signal, which allows for their better detection.
Since the imputed values for the seven CSSA models described above provide approximately equal errors on the artificial gaps, the average of these values can be recommended as the final method of filling in the actual gap in the IERS C01 series. This filling method using the average of seven gap fillings has a slightly smaller MAE (0.0628 vs. 0.0638–0.0662) on the considered set of artificial gaps, compared to each of these seven gap fillings.
5 Final results
Average results (filling $Y_{p}$ values) computed using the BCA and SSA methods are presented in Table 4 with the number of decimal places corresponding to the IERS C01 series. The IERS C01 series with added filling values is shown in Fig. 7.
Table 4: Filled $Y_{p}$ values, arcseconds.
Epoch Model
BCA SSA
1858.9 $-0.236505$ $-0.247675$
1859.0 $-0.153508$ $-0.164798$
1859.1 $-0.050599$ $-0.053550$
1859.2 $0.039868$ $0.045943$
1859.3 $0.091581$ $0.100224$
1859.4 $0.092240$ $0.095042$
1859.5 $0.046379$ $0.038934$
1859.6 $-0.027558$ $-0.042570$
1859.7 $-0.104471$ $-0.118523$
1859.8 $-0.161363$ $-0.164565$
1859.9 $-0.184487$ $-0.170918$
1860.0 $-0.172543$ $-0.143915$
1860.1 $-0.135271$ $-0.101195$
1860.2 $-0.088418$ $-0.063027$
1860.3 $-0.047375$ $-0.043324$
1860.4 $-0.021975$ $-0.044098$
1860.5 $-0.014260$ $-0.055752$
1860.6 $-0.019681$ $-0.063258$
1860.7 $-0.030755$ $-0.055410$
1860.8 $-0.041278$ $-0.032369$
1860.9 $-0.049166$ $-0.006942$
Both methods show good agreement, especially in the first year of the gap interval, 1859–1860. The agreement in the second year of the interval, 1860–1861, is somewhat worse, but can be considered quite satisfactory, considering that the $Y_{p}$ error in the IERS C01 series is 0.09′′. Generally speaking, it should be kept in mind that the accuracy of the reconstruction is limited by the accuracy of the IERS C01 data, especially before the gap in the 1840s–1850s.
6 Conclusions
In this paper, the first attempt has been made to address the problem of filling the 2-year gap in the IERS C01 $Y_{p}$ series in 1858.9–1860.9. For this purpose, two methods were proposed and investigated. The first is the LS adjustment of the 9-parameter BCA model, consisting of a bias and two harmonic terms, CW and AW, with linearly changing amplitude, and its modification, the 13-parameter BCA2 model, which includes two CW components. The second method is CSSA (complex SSA), applied to the IERS C01 PM series in complex form, which allows us to obtain a fully data-driven solution, not limited by the parameters of parametric models. Besides, the discrepancies between the results obtained by the two methods may be due to the fact that the parametric models are based on a several times shorter interval than the SSA model. Thus, parametric models mainly describe the local PM behaviour on the date interval around the gap and under the assumption of a linear change in the CW and AW amplitudes, while the SSA model is much less sensitive to a priori assumptions and is more suitable for describing the long-term PM features, which is the primary interest of this paper.
Therefore, we consider the result obtained by SSA as more reliable because it takes into account all the features of the observed PM data and all the main components of its structure, while the parametric model uses only predefined components, which may not adequately describe the results of the observations. However, we present both the results obtained by the BCA and SSA methods in Table 4 so that the user can choose the variant that looks most appropriate for the problem under investigation, or simply use the average of the two.
It should be noted that there are many methods for efficiently processing unevenly spaced data, in particular, data with gaps. Many of widely used methods of data analysis, such as LS adjustment, spectral analysis, SSA, wavelet technique, etc., have modifications for processing unevenly spaced data. Many examples of application of such analyses to study PM series can be found in the literature, e. g., Gaposchkin (1972); Kołaczek and Kosek (1999); Gambis (2000); Guo et al. (2005); Zotov and Bizouard (2012); Beutler et al. (2020); Lopes et al. (2022). However, the missing two years of $Y_{p}$ data may introduce bias in PM modeling if inappropriate analysis methods are used. Our work aims to provide interested users with the option to work with evenly distributed data and to use appropriate mathematical methods, if they want. In most cases, one cannot expect a significant difference in the results obtained with original and filled series, but working with evenly spaced data is more convenient and allows for detailed epoch-by-epoch analysis. Including the proposed filling values into the IERS C01 series (indeed, with corresponding note) will provide full continuous, evenly spaced PM series over the entire date range starting from 1846. This will also allow to avoid (even small) mistakes in using this series for scientific analysis if the gap in the series in its current, formally evenly spaced form is not properly taken into account. Perhaps for this reason, e. g., Schuh et al. (2001) and Vondrák (1988) started their analysis from 1861.0.
In this study, mathematical methods were mainly used to complete the existing IERS C01 PM series by filling in the missing $Y_{p}$ values for 21 epochs in the period 1858.9 to 1860.9. On the other hand, mining and reprocessing historical observations looks very interesting, but this huge task goes far beyond the scope of this paper and will obviously require cooperation between observatories that could preserve in their archives declination and latitude observations carried out in the 19th century. Therefore, attempts to improve the quality of the IERS C01 series in the 19th century should be considered, in our opinion, an actual task.
Appendix A Filling the gaps with SSA
@@
Suppose there is a method that knows how to remove noise from a noisy signal. Then we can propose a simple iterative gap-filling method. First, the places with missing values are filled with some initial values, e. g., the mean of the series, then the signal extraction is performed, the obtained signal estimates are put at the gaps, the signal extraction method is applied again, and so the procedure is repeated until convergence.
@@
In this paper, we use SSA (Golyandina et al., 2001; Golyandina and Zhigljavsky, 2020) as the signal reconstruction (noise removal) method that is part of the iterative gap-filling process (Kondrashov and Ghil, 2006). Let us describe it briefly. The SSA method in its version for signal extraction has only two parameters, the window length $L$ and the number of components $r$. At the first stage, the time series $\mathsf{X}=(x_{1},\ldots,x_{N})$ is transformed into the so-called trajectory matrix:
@@
using the parameter $L$. Next, the singular value decomposition (SVD) of the trajectory matrix is constructed. Let $\mathbf{S}=\mathbf{X}\mathbf{X}^{\mathrm{T}}$, $\lambda_{1},\ldots,\lambda_{L}$ be the eigenvalues of the matrix $\mathbf{S}$ taken in non-increasing order, $U_{1},\ldots,U_{L}$ be the orthonormalized system of eigenvectors corresponding to these eigenvalues. Let $d=\max\{k:\lambda_{k}>0\}$ and $V_{k}=\mathbf{X}^{\mathrm{T}}U_{k}/\sqrt{\lambda_{k}}$, $k=1,\ldots,d$. Then the SVD of the matrix $\mathbf{X}$ can be represented as a sum of elementary matrices:
@@
In the second step, the first $r$ elementary matrices of the SVD are summed:
@@
and their sum is converted into a time series
@@
The SSA algorithm can be extended to the simultaneous analysis of two time series, $\mathsf{X}$ and $\mathsf{Y}$, in two forms. In the case of Complex SSA (CSSA) for a time series of the form $x_{n}+{\mathrm{i}\mkern 1.0mu}y_{n}$, $n=1,\ldots,N$, the transpose operator $\mathrm{T}$ is replaced by the Hermite conjugate operator $\mathrm{H}$. The complex form of SSA has been considered since its origin. In the complex form, the first descriptions of the method can be found, for example, in Kumaresan and Tufts (1982); Keppenne and Lall (1996).
In the case of Multivariate SSA (MSSA) for two time series of the $(x_{n},y_{n})$, $n=1,\ldots,N$, the difference lies in the construction of the trajectory matrix, which consists of stacked trajectory matrices of the time series. A more detailed description can be found in Golyandina et al. (2015).